# COMPASS: Computational Mapping of Patient-Therapist Alliance Strategies with Language Modeling

### v1

@@ 4.1 Working Alliance Analysis
The analysis of past therapy sessions, as well as real-time sessions, has the potential to help trained therapists identify key segments of therapy leading to breakthroughs. By leveraging computational modeling and statistical optimization, therapists can compound their expertise with causal and predictive analytic modeling to enhance their understanding of the therapeutic process. Additionally, trainees can benefit from studying annotated versions of sessions conducted by experts, further sharpening their intuition and skills. Furthermore, the integration of generative language models and statistical optimization techniques may enable the design of limited chatbots for triage and emergency response in mental health care [30].
@@
Furthermore, our research faces the challenge of limited clinical validation in the field. Working alliance, as a proxy for therapeutic alignment and outcome, relies on approximate measures and operational methods. Existing approaches provide approximations to psychological properties and lack a gold standard. In future work, clinical studies are needed to validate our results in real-world settings and intervention contexts.
@@ 4.2 Topic Modeling and Interpretability
The analysis of the temporal dynamics within the learned topics provides further insights into the therapeutic process. By mapping the topic scores of dialogue turns to the working alliance scores, we can identify topics and dialogue segments that are potentially indicative of therapeutic breakthroughs. For instance, in depression sessions, the topic related to self-esteem may be associated with improvements in the bond and task scales of the working alliance. Similarly, in schizophrenia sessions, the topic related to family dynamics may contribute to positive changes in the bond scale.
@@
In summary, our analysis of the working alliance and topic modeling in psychotherapy transcripts offers valuable insights into the therapeutic relationship and the content of therapy sessions. These findings contribute to our understanding of the therapeutic process and have implications for improving mental health care. While there are limitations and challenges in inferring psychological states from text data and validating our approach clinically, future studies and advancements in the field can help address these issues and further enhance the effectiveness of AI in mental health care.
Additional Notes to Cautious and Responsible Interpretations. It is crucial to acknowledge that these summaries are based on the specific patient populations represented in each dataset and may not capture the full range of therapeutic interventions employed by therapists. Each dataset reflects unique challenges and therapeutic goals associated with the corresponding patient population, necessitating tailored approaches and interventions by the therapists.
While these results are informative to a certain degree, we would like to acknowledge the limitations to the methodologies presented. The data we used to train the topic models do not necessarily represent different clinical conditions in a balanced way. In the classification validation task, we have imposed iteratively sampling to tackle the imbalance issue, but for training the topic models, we used the full text corpus available, as certain clinical labels (e.g. suicidal patients) only have a handful of sessions available (e.g. 11). Due to the limited access to therapy sessions with suicidal patents, the topics characterized by the topic models might be less representative to the conversations happening in this particular patient groups. The Alexander Street dataset mentions that they have more clinical conditions other than the analyzed 4 classes, but due to the licensing and access limitations, we can only obtain the 4 classes we presented. As open science and data sharing initiatives in the psychiatry domains become more prominent, we believe our methodologies can be adapted in a responsible way to a broader spectrum of clinical conditions.
@@ 4.3 Ethical Considerations
First and foremost, it is important to emphasize that our intention is not to replace or diminish the role of psychiatrists or therapists, but rather to assist them in their practice. The goal is to provide valuable insights and support to both experienced therapists and junior psychiatrists, particularly in the educational setting, where the interpretability of our models can inform the strategies employed by seasoned therapists and aid in the training of future mental health professionals.
When working with patient data, privacy and security are of utmost importance. We have followed ethical guidelines and operational suggestions [32, 33, 34] to ensure the proper anonymization and protection of sensitive information. The dataset we analyzed was sourced from ProQuest’s Alexander Street platform, and all personally identifiable information, such as metadata, user names, identifiers, and doctors’ names, has been removed.
@@
In the context of mental health and psychological well-being, there are additional ethical considerations. The emergence of wearable devices, digital health records, brain imaging measurements, smartphone applications, and social media has transformed the landscape of monitoring and treating mental health conditions. However, it is important to approach these advances with caution, as many of these technologies are still in the proof-of-concept stage [34]. Rigorous clinical validation and regulatory approval are necessary before deploying these technologies for patient care and therapeutic decision-making.
It is crucial to acknowledge the limitations of our work and the challenges in the field. Machine learning solutions in psychiatry, including our approach, face difficulties in conducting systematic clinical validation and ensuring the generalizability of results [35]. Real-world applications often involve small sample sizes, missing data points, and highly correlated variables, which can impact the generalizability and reliability of machine learning models. Therefore, it is essential to exercise caution when interpreting and applying the results of such models.
@@
To ensure responsible and safe deployment of AI systems in mental health care, it is necessary to be mindful of potential biases and ethical challenges. Gender bias, language-related ambiguity, and ethnicity-related mental illness connections are examples of such challenges [36]. Practitioners and machine learning researchers must be aware of these issues and take steps to mitigate biases and promote fairness in AI systems. Engaging the public in discussions about the usage of AI in mental health care is important to foster awareness and avoid unrealistic expectations of AI as a “domain expert” [37].
In our analysis, we utilized a dataset of over 950 psychotherapy transcripts. While it is the largest dataset available in this research domain, we acknowledge the limitations in terms of its representativeness and generalizability to all populations. The anonymized nature of the dataset and the lack of detailed information about the collection process and demographics of the participants pose constraints. However, we believe that the insights gained from our interpretable investigations are unlikely to increase unforeseeable risks to the patients and have the potential to be valuable in clinical practice.
We are committed to upholding ethical standards in our work. We prioritize patient privacy and data security, acknowledge the limitations and challenges in the field, and strive to ensure responsible and unbiased deployment of AI systems in mental health care. By addressing these ethical considerations, we aim to contribute to the advancement of AI technologies in a manner that benefits patients, clinicians, and the field of mental health as a whole.
@@ 5 Conclusions
Furthermore, we have incorporated temporal modeling into topic modeling to parse topics in different segments of the therapy sessions. This temporal analysis adds another layer of interpretability and enables us to observe session trajectories and their separability between patients and therapists. We have noted that in anxiety and depression sessions, the trajectories of patients and therapists tend to be more separable, whereas in schizophrenia sessions, they are more entangled. This initial step toward turn-level resolution temporal analysis in topic modeling provides valuable insights that can help therapists improve the effectiveness of psychotherapy.
@@ Acknowledgments
The authors thank Ravi Tejwani, Barnaby Nelson and Alison Yung for helpful discussions and suggestions.
Supplementary Material
We compare the semantic similarity of the working alliance inventories with the transcripts using Algorithm 1. The dialogue between the patient and therapist is transcribed into pairs of turns denoted as for the patient’s response turn, followed by the therapist’s response turn . The inventories of working alliance questionnaires are also provided in pairs: for the patient and for the therapist, each comprising 36 statements. We use sentence or paragraph embeddings to encode both the dialogue turns and the inventories, and then compute the cosine similarity between the embedding vectors of each turn and its corresponding inventory vectors. This approach yields a 36-dimensional working alliance score for each turn.
For classification tasks, we concatenate the 36-dimensional working alliance scores estimated from the current turn with the sentence embedding of the turn. This combined feature vector is then fed into the Working Alliance Transformer (WAT) and Working Alliance LSTM (WA-LSTM) Algorithm 2, which are based on the Transformer architecture [24] and the LSTM network model [25] respectively. The WAT and WA-LSTM serve as sequence classifiers, taking in the sequence of feature vectors and predicting the clinical condition associated with the sequence.
@@
The pipeline for temporal topic modeling analysis (TMM) is outlined in Algorithm 3. For each turn in the transcript, we calculate a topic score vector with each dimension representing the likelihood of the turn belonging to a specific topic. To characterize the directional property of each turn with a particular topic, we compute the cosine similarity between the embedded topic vector and the embedded turn vector. This approach provides a measure of the topic relevance to each turn in the sequence.
| Anxiety | Depression | Schizophrenia | ||||
| Topic diversity | Topic coherence | Topic diversity | Topic coherence | Topic diversity | Topic coherence | |
| NVDM-GSM | 0.653 | -380.933 | 0.487 | -316.439 | 0.527 | -431.393 |
| WTM-MMD | 0.927 | -453.929 | 0.907 | -359.964 | 0.447 | -403.694 |
| WTM-GMM | 0.907 | -425.515 | 0.340 | -236.815 | 0.467 | -204.930 |
| ETM | 0.893 | -449.000 | 0.933 | -367.069 | 0.973 | -310.211 |
| BATM | 0.720 | -441.049 | 0.773 | -443.394 | 0.500 | -337.825 |
| Anxiety | Depression | Schizophrenia | ||||||||||
| NVDM-GSM | 0.410 | 0.484 | -0.844 | -0.019 | 0.495 | 0.531 | -3.522 | -0.109 | 0.642 | - | -1.954 | -0.065 |
| WTM-MMD | 0.340 | 0.428 | -2.827 | -0.099 | 0.290 | 0.462 | -3.797 | -0.124 | 0.576 | 0.751 | -0.997 | -0.036 |
| WTM-GMM | 0.353 | 0.413 | -3.259 | -0.116 | 0.678 | 0.535 | -0.126 | -0.006 | 0.572 | 0.774 | -1.587 | -0.050 |
| ETM | 0.413 | - | -2.903 | -0.093 | 0.403 | - | -2.399 | -0.05 | 0.379 | 0.864 | -7.232 | -0.199 |
| BATM | 0.352 | 0.387 | -5.056 | -0.190 | 0.404 | 0.423 | -4.238 | -0.160 | 0.507 | 0.816 | -9.655 | -0.343 |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 1.600e-22 | - | ||
| schizophrenia | 2.998e-05 | 3.249e-01 | - | |
| suicidal | 6.004e-01 | 2.176e-01 | 3.908e-01 | - |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 2.207e-13 | - | ||
| schizophrenia | 1.565e-11 | 3.883e-03 | - | |
| suicidal | 3.627e-23 | 2.555e-29 | 3.629e-34 | - |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 7.186e-08 | - | ||
| schizophrenia | 3.010e-02 | 4.959e-01 | - | |
| suicidal | 1.115e-10 | 3.532e-13 | 3.249e-12 | - |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 1.177e-01 | - | ||
| schizophrenia | 8.006e-04 | 4.519e-05 | - | |
| suicidal | 2.148e-01 | 1.289e-01 | 9.510e-01 | - |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 7.186e-08 | - | ||
| schizophrenia | 3.010e-02 | 4.959e-01 | - | |
| suicidal | 1.115e-10 | 3.532e-13 | 3.249e-12 | - |
| anxiety | depression | schizophrenia | suicidal | |
@@
| depression | 1.177e-01 | - | ||
| schizophrenia | 8.006e-04 | 4.519e-05 | - | |
| suicidal | 2.148e-01 | 1.289e-01 | 9.510e-01 | - |

### v2

@@ 4.1 Working Alliance Analysis
The analysis of past therapy sessions, as well as real-time sessions, has the potential to help therapists identify key segments of therapy leading to breakthroughs. By leveraging computational modeling and statistical optimization, therapists can compound their expertise with causal and predictive analytic modeling to enhance their understanding of the therapeutic process. Additionally, trainees can benefit from studying annotated versions of sessions conducted by experts, further sharpening their intuition and skills. Furthermore, the integration of generative language models and statistical optimization techniques may enable the design of limited chatbots for triage and emergency response in mental health care [40].
@@
Furthermore, our research faces the challenge of limited clinical validation in this field. While the Working Alliance Inventory (WAI) is a widely used measure for assessing working alliance, it is not a standard practice across all therapeutic settings and does not represent a comprehensive gold standard, particularly for granular, turn-by-turn analyses. Current methods rely on approximate measures and operational approaches that cannot fully capture the nuanced, real-time dynamics of therapeutic interactions. In future work, clinical studies incorporating detailed, turn-by-turn evaluations by human annotators, albeit challenging and resource-intensive, may provide more direct validation of inferred alliance scores within real-world settings and intervention contexts.
Lastly, it is important to note that the dataset had limited access to sessions involving suicidal patients, which might impact the robustness of our findings related to this group.
@@ 4.2 Topic Modeling and Interpretability
The analysis of the temporal dynamics within the learned topics provides further insights into the therapeutic process. By mapping the topic scores of dialogue turns to the working alliance scores, we can identify topics and dialogue segments that are potentially indicative of therapeutic breakthroughs. For instance, in depression sessions, the topic related to self-esteem may be associated with improvements in the bond and task scales of the working alliance. Similarly, in sessions with patients with schizophrenia, the topic related to family dynamics may contribute to positive changes in the bond scale.
@@
While these results are informative to a certain degree, we would like to acknowledge the limitations to the methodologies presented. The data we used to train the topic models do not necessarily represent different clinical conditions in a balanced way. In the classification validation task, we have imposed iterative sampling to tackle the imbalance issue, but for training the topic models, we used the full text corpus available, as certain clinical labels (e.g. suicidal patients) only have a handful of sessions available. Due to the limited access to therapy sessions with suicidal patents, the topics characterized by the topic models might be less representative to the conversations happening in this particular patient groups. The Alexander Street dataset mentions that they have more clinical conditions other than the analyzed 4 classes, but due to access limitations, we can only obtain the 4 classes we presented. As open science and data sharing initiatives in the psychiatry domains become more prominent, we believe our methodologies can be adapted in a responsible way to a broader spectrum of clinical conditions.
@@ 4.3 Ethical Considerations
First and foremost, our approach was designed fundamentally as a companion tool to provide valuable insights and support to both experienced therapists and junior clinicians, particularly in the educational setting, where the interpretability of our models can inform the strategies employed by seasoned therapists and aid in the training of future mental health professionals.
@@
When working with patient data, privacy and security are of utmost importance. We have followed ethical guidelines and operational suggestions [42, 43, 44] to ensure the proper anonymization and protection of sensitive information. The dataset we analyzed had all personally identifiable information, such as metadata, user names, identifiers, and doctors’ names, throughly removed. In the context of mental health and psychological well-being, there are additional ethical considerations. The emergence of wearable devices, digital health records, brain imaging measurements, smartphone applications, and social media has transformed the landscape of monitoring and treating mental health conditions. However, it is important to approach these advances with caution, as many of these technologies are still in the proof-of-concept stage [44]. Rigorous clinical validation and regulatory approval are necessary before deploying these technologies for patient care and therapeutic decision-making. Machine learning solutions in psychiatry, including our approach, face difficulties in conducting systematic clinical validation and ensuring the generalizability of results [45]. Real-world applications often involve small sample sizes, missing data points, and highly correlated variables, which can impact the generalizability and reliability of machine learning models.
@@
To ensure responsible and safe deployment of AI systems in mental health care, it is necessary to be mindful of potential biases and ethical challenges. Gender bias, language-related ambiguity, and ethnicity-related mental illness connections are examples of such challenges [46]. Practitioners and machine learning researchers must be aware of these issues and take steps to mitigate biases and promote fairness in AI systems. Engaging the public in discussions about the usage of AI in mental health care is important to foster awareness and avoid unrealistic expectations of AI as a “domain expert” [47]. We utilized a dataset of over 950 psychotherapy transcripts. While it is the largest dataset available in this research domain, we acknowledge the limitations in terms of its representativeness and generalizability to all populations. The anonymized nature of the dataset and the lack of detailed information about the collection process and demographics of the participants pose constraints. However, we believe that the insights gained from our interpretable investigations are unlikely to increase unforeseeable risks to the patients and have the potential to be valuable in clinical practice.
@@ 5 Conclusions
Furthermore, we have incorporated temporal modeling into topic modeling to parse topics in different segments of the therapy sessions. This temporal analysis adds another layer of interpretability and enables us to observe session trajectories and their separability between patients and therapists. We have noted that in sessions with patients with anxiety and depression, the trajectories of patients and therapists tend to be more separable, whereas in schizophrenia sessions, they are more entangled. This initial step toward turn-level resolution temporal analysis in topic modeling provides valuable insights that can help therapists improve the effectiveness of psychotherapy.
@@ Acknowledgments
The authors thank Drs. Barnaby Nelson, Alison Yung, and Ravi Tejwani for helpful discussions and suggestions.
Author Contributions
BL initiated the study, designed the research, conducted the analysis, wrote the original manuscript, and obtained funding. GC contributed to research design and provided guidance on manuscript writing. CC, YL, and DB provided advice on the study and manuscript writing. All authors reviewed, edited, contributed to, and approved the final manuscript.
Supplementary Materials
We compare the semantic similarity of the working alliance inventories with the transcripts using Algorithm S1. The dialogue between the patient and therapist is transcribed into pairs of turns denoted as for the patient’s response turn, followed by the therapist’s response turn . The inventories of working alliance questionnaires are also provided in pairs: for the patient and for the therapist, each comprising 36 statements. We use sentence or paragraph embeddings to encode both the dialogue turns and the inventories, and then compute the cosine similarity between the embedding vectors of each turn and its corresponding inventory vectors. This approach yields a 36-dimensional working alliance score for each turn.
For classification tasks, we concatenate the 36-dimensional working alliance scores estimated from the current turn with the sentence embedding of the turn. This combined feature vector is then fed into the Working Alliance Transformer (WAT) and Working Alliance LSTM (WA-LSTM) Algorithm S2, which are based on the Transformer architecture [34] and the LSTM network model [35] respectively. The WAT and WA-LSTM serve as sequence classifiers, taking in the sequence of feature vectors and predicting the clinical condition associated with the sequence.
@@
The pipeline for temporal topic modeling analysis (TMM) is outlined in Algorithm S3. For each turn in the transcript, we calculate a topic score vector with each dimension representing the likelihood of the turn belonging to a specific topic. To characterize the directional property of each turn with a particular topic, we compute the cosine similarity between the embedded topic vector and the embedded turn vector. This approach provides a measure of the topic relevance to each turn in the sequence.
Supplementary Tables
| Scale | Inventory item | Directionality |
| BOND | I felt uncomfortable with _. | - |
| TASK | _ and I agreed about the things I will need to do in therapy to help improve my situation. | + |
| GOAL | I was worried about the outcome of the sessions. | - |
| TASK | What I was doing in therapy gave me new ways of looking at my problem. | + |
| BOND | _ and I understood each other. | + |
| GOAL | _ perceived accurately what my goals were. | + |
| TASK | I find what I was doing in therapy confusing. | - |
| BOND | I believe _ liked me. | + |
| GOAL | I wish _ and I could have clarified the purpose of our sessions. | - |
| GOAL | I disagreed with _ about what I ought to get out of therapy. | - |
| TASK | I believe the time _ and I were spending together was not spent efficiently. | - |
| GOAL | _ did not understand what I was trying to accomplish in therapy. | - |
| TASK | I was clear on what my responsibilities were in therapy. | + |
| GOAL | The goals of the sessions were important for me. | + |
| TASK | I find what _ and I were doing in therapy was unrelated to my concerns. | - |
| TASK | I feel that the things I did in therapy helped me to accomplish the changes that I wanted. | + |
| BOND | I believe _ was genuinely concerned for my welfare. | + |
| TASK | I was clear as to what _ wanted me to do in those sessions. | + |
| BOND | _ and I respected each other. | + |
| BOND | I feel that _ was not totally honest about his/her feelings toward me. | - |
| BOND | I was confident in _’s ability to help me. | + |
| GOAL | _ and I were working towards mutually agreed upon goals. | + |
| BOND | I feel that _ appreciated me. | + |
| TASK | We agreed on what was important for me to work on. | + |
| GOAL | As a result of the therapy I became clearer as to how I might be able to change. | + |
| BOND | _ and I trusted one another. | + |
| GOAL | _ and I had different ideas on what my problems were. | - |
| BOND | My relationship with _ was very important to me. | + |
| BOND | I had the feeling that if I said or did the wrong things, _ would stop working with me. | - |
| GOAL | _ and I collaborated on setting goals for my therapy. | + |
| TASK | I was frustrated by the things I was doing in therapy. | - |
| GOAL | We had a good understanding of the kind of changes that would be good for me. | + |
| TASK | The things that _ was asking me to do did not make sense. | - |
| GOAL | I did not know what to expect as the result of my therapy. | - |
| TASK | I believe the way we were working with my problem was correct. | + |
| BOND | I feel _ cared about me even when I did things that he/she did not approve of. | + |
| Anxiety | Depression | Schizophrenia | ||||||
| Topic diversity | Topic coherence | Topic diversity | Topic coherence | Topic diversity | Topic coherence | |||
| NVDM-GSM | 0.653 | -380.933 | 0.487 | -316.439 | 0.527 | -431.393 | ||
| WTM-MMD | 0.927 | -453.929 | 0.907 | -359.964 | 0.447 | -403.694 | ||
| WTM-GMM | 0.907 | -425.515 | 0.340 | -236.815 | 0.467 | -204.930 | ||
| ETM | 0.893 | -449.000 | 0.933 | -367.069 | 0.973 | -310.211 | ||
| BATM | 0.720 | -441.049 | 0.773 | -443.394 | 0.500 | -337.825 | ||
| Anxiety | Depression | Schizophrenia | ||||||||||||
| NVDM-GSM | 0.410 | 0.484 | -0.844 | -0.019 | 0.495 | 0.531 | -3.522 | -0.109 | 0.642 | - | -1.954 | -0.065 | ||
| WTM-MMD | 0.340 | 0.428 | -2.827 | -0.099 | 0.290 | 0.462 | -3.797 | -0.124 | 0.576 | 0.751 | -0.997 | -0.036 | ||
| WTM-GMM | 0.353 | 0.413 | -3.259 | -0.116 | 0.678 | 0.535 | -0.126 | -0.006 | 0.572 | 0.774 | -1.587 | -0.050 | ||
| ETM | 0.413 | - | -2.903 | -0.093 | 0.403 | - | -2.399 | -0.05 | 0.379 | 0.864 | -7.232 | -0.199 | ||
| BATM | 0.352 | 0.387 | -5.056 | -0.190 | 0.404 | 0.423 | -4.238 | -0.160 | 0.507 | 0.816 | -9.655 | -0.343 | ||
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | **** (1.550e-26) | - | ||
| Schizophrenia | *** (3.835e-05) | ns (1.267e-01) | - | |
| Suicidality | ns (6.521e-01) | * (1.900e-02) | ns (6.568e-02) | - |
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | **** (1.796e-25) | - | ||
| Schizophrenia | ** (1.937e-03) | **** (7.263e-17) | - | |
| Suicidality | **** (8.197e-20) | **** (2.108e-28) | **** (5.465e-14) | - |
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | **** (1.600e-22) | - | ||
| Schizophrenia | *** (2.998e-05) | ns (3.249e-01) | - | |
| Suicidality | ns (6.004e-01) | ns (2.176e-01) | ns (3.908e-01) | - |
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | **** (2.207e-13) | - | ||
| Schizophrenia | **** (1.565e-11) | ** (3.883e-03) | - | |
| Suicidality | **** (3.627e-23) | **** (2.555e-29) | **** (3.629e-34) | - |
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | **** (7.186e-08) | - | ||
| Schizophrenia | * (3.010e-02) | ns (4.959e-01) | - | |
| Suicidality | **** (1.115e-10) | **** (3.532e-13) | **** (3.249e-12) | - |
| Anxiety | Depression | Schizophrenia | Suicidality | |
@@
| Depression | ns (1.177e-01) | - | ||
| Schizophrenia | *** (8.006e-04) | **** (4.519e-05) | - | |
| Suicidality | ns (2.148e-01) | ns (1.289e-01) | ns (9.510e-01) | - |